# WildLive: Near Real-time Visual Wildlife Tracking onboard UAVs

### v2

@@ 7 Conclusion and Future Work
When collecting our WildLive Benchmark Dataset and during test flights, we adhered to strict ethical standards to ensure the wellbeing of the animals and the appropriateness and integrity of the research. Our data was collected using drones in the Ol Pejeta Conservancy, Kenya. Drones were operated at safe distances to avoid causing stress or disturbance to animals. Additionally, all recordings were made in a non-invasive manner, with no direct interaction with wildlife. To further protect privacy and adhere to ethical guidelines, we ensured that human faces are not contained in recordings, focusing solely on animals and their behavior. All data collection procedures followed relevant local regulations, such as those provided by the Kenya Civil Aviation Authority (KCAA), the Kenya Wildlife Research and Training Institute (KWRTI), and other ethical guidelines. Efforts were made to respect the natural behaviours of the animals maximally. By prioritising animal welfare, privacy, and ethical research practices, the dataset can contribute to scientific advancements while minimising any harm to the environment and its inhabitants.

### v3

@@ Acknowledgements
When collecting our WildLive Benchmark Dataset and during test flights, we adhered to strict ethical standards to ensure the wellbeing of the animals and the appropriateness and integrity of the research. Our data was collected using drones in the Ol Pejeta Conservancy, Kenya. Drones were operated at safe distances to avoid causing stress or disturbance to animals. Additionally, all recordings were made in a non-invasive manner, with no direct interaction with wildlife. To further protect privacy and adhere to ethical guidelines, we ensured that human faces are not contained in recordings, focusing solely on animals and their behavior. All data collection procedures followed relevant local regulations, such as those provided by the Kenya Civil Aviation Authority (KCAA)(Authorization number KCAA/UAS/OPS/0068/2025), the Kenya Wildlife Research and Training Institute (KWRTI), and other ethical guidelines. Efforts were made to respect the natural behaviours of the animals maximally. By prioritising animal welfare, privacy, and ethical research practices, the dataset can contribute to scientific advancements while minimising any harm to the environment and its inhabitants.